# Split Unlearning

### v3

@@ Abstract
We for the first time propose, implement, and evaluate a practical Split Unlearning framework by enabling SISA-based machine unlearning (SP’21) in Split Learning (SL).
We introduce SplitWiper and SplitWiper+, which leverage the inherent “Sharded” structure of SL and address the issues where existing SL methods compromise the “Isolated” principle of SISA due to the tight coupling between clients and the server. This coupling involves frequent bidirectional data flows and iterative training across all clients. We resolve these challenges by designing a new one-way-one-off propagation scheme. Our design decouples the propagation of neural signals between clients and between clients and the server, enabling SISA-based unlearning in SplitWiper, even with absent clients. SplitWiper+ further enhances client label privacy against the server under this scheme by using differential privacy. Our extensive experiments across diverse data distributions and tasks demonstrate that SplitWiper, which involves only the requesting clients, achieves 0% unlearning accuracy and is over 108% more effective in retaining accuracy than existing SL methods, while maintaining constant overhead through its one-way-one-off propagation, reducing computational and communication costs by over 99%. SplitWiper+ preserves over 90% of label privacy when sharing masked labels with servers.
@@ 1 Introduction
Split Learning (SL) is a key approach in privacy-preserving distributed learning, particularly in scenarios where clients with limited computational resources, such as smartphones or IoT devices, need to offload part of the training computations to servers [1, 2, 3]. Existing SL frameworks primarily include two configurations: Vanilla SL and U-shaped SL [4]. In Vanilla SL, the model is split into two segments: clients manage the segment receiving the training data and send intermediate values to the server, while the server handles the segment producing the final outputs without accessing raw data. U-shaped SL extends this by eliminating the need for label sharing, protecting both label privacy and data privacy. The model in U-shaped SL is divided into three segments: clients process the input data and produce output classes, while the server manages the intermediate layers.
@@
Ensuring compliance with regulatory requirements for “data to be forgotten” (e.g., GDPR [5]), which grants individuals the right to delete their personal data, necessitates the development of machine unlearning [6] techniques to effectively remove specific data from models. However, the requirement becomes challenging within the context of SL, where the model segments across clients and servers need to be jointly updated. Current methods include retraining the model [7, 8] or using alternatives like fine-tuning, gradient ascent, and knowledge distillation [9, 10, 11]. While these alternatives are favored for their lower unlearning efficiency (i.e. time consumption) compared to retraining, which prioritizes effectiveness (i.e., accuracy[12], they all require the inclusion of all data samples during each update iteration. This becomes problematic when some data samples are absent due to contextual constraints. The Sharded, Isolated, Sliced, and Aggregated (SISA) method [6, 13, 14] addresses it by segmenting data and training multiple models in parallel, maintaining higher effectiveness and efficiency than non-SISA methods.
@@
Research gaps. Existing SL frameworks [1, 2, 3] are inadequate for directly applying SISA in multi-client scenarios, where multiple clients have their independent models while sharing a common model on the server. Although the clients’ training data can be considered “Sharded”, SL undermines the “Isolated” requirement of SISA for two reasons: (1) the server model is tightly coupled with the client models through frequent bidirectional propagation, and (2) the shared server model is iteratively trained across all clients. Therefore, a new framework is necessary to decouple the shared server model from client models, enabling SISA-based unlearning in multi-client SL environments.
@@
Contributions. In this paper, we address these gaps in enabling SISA-based unlearning in multi-client SL environments where unlearning can thus be conducted smoothly even with absent clients. To be specific:
@@ 2.1 Split Learning
Key factors. SL is a distributed machine learning technique designed to train neural networks without requiring the sharing of raw data between $[1,K]$ clients, each owning datasets $D_{o}^{1},D_{o}^{2},\ldots,D_{o}^{K}$, and servers that hold no data [4, 15]. A typical SL solution involves two key factors: topology and training.
@@
    Topology. The network is divided into two parts, referred to as client-side and server-side. We denote the network as a sequence of layers $\{L_{1},L_{2},\ldots,L_{n}\}$. The topology step involves splitting this sequence into two subsets: the client-side layers for $K$ clients in parallel are denoted as $\{\{L_{1}^{1},L_{2}^{1},\ldots,L_{m}^{1}\}$, $\ldots$, $\{L_{1}^{K},L_{2}^{K},\ldots,L_{m}^{K}\}\}$ and the server-side layers $\{L_{m+1},L_{m+2},\ldots,L_{n}\}$. The clients and server initialize their respective portions of the network randomly.
@@
    Training. Each client-$k$ processes its local data through the initial layers of the model, producing an intermediate output $H_{k}^{(t)}=\mathcal{F}_{o}^{k}(X_{k};W_{k}^{(t)})$ through its layers $\{L_{1}^{k},L_{2}^{k},\ldots,L_{m}^{k}\}$, where $W_{k}^{(t)}$ represents client-$k$’s weights at iteration $t$. This output is sent to the server, which continues the forward propagation through its layers $\{L_{m+1},L_{m+2},\ldots,L_{n}\}$ and updates its own weights $W_{s}^{(t)}$ based on the gradients: $W_{s}^{(t+1)}=W_{s}^{(t)}-\eta\Phi\left(\frac{\partial\mathcal{L}_{1}}{\partial W_{s}^{(t)}},\frac{\partial\mathcal{L}_{2}}{\partial W_{s}^{(t)}},\dots,\frac{\partial\mathcal{L}_{K}}{\partial W_{s}^{(t)}}\right)$, where $\frac{\partial\mathcal{L}_{k}}{\partial W_{s}^{(t)}}$ is the gradient contribution of loss values from client-$k$ and $\Phi(\cdot)$ is a customized aggregation function. The server then sends back the gradients to each client, allowing them to update their own weights: $W_{k}^{(t+1)}=W_{k}^{(t)}-\eta\frac{\partial\mathcal{L}}{\partial W_{k}^{(t)}}$.
The processes described above enable training without the need to share raw data or specific model details between the client and server. In scenarios involving multiple clients and a single server, two primary training methods are commonly employed. The first method synchronizes model weights between clients after each epoch, while the second method operates without synchronization, with clients alternating epochs in collaboration with the server [4].
Incremental training process. This process has been utilized by many prevalent learning frameworks including Federated Learning (FL) [16, 17] and SL, where a model is continuously updated by gradually incorporating new data over time. Each iteration of back-propagation needs to be sent back to every client involved in the training process and builds incrementally on the previous ones, embodying the incremental nature of SL [18]. This process enables the model to dynamically learn and adapt as new data becomes available, ensuring it remains responsive to evolving patterns and trends.
The challenges of applying SISA in multi-client SL scenarios stem from the incremental training process inherent in these setups. This iterative process across multiple clients undermines the Isolated requirement of SISA due to the close integration of the server and client model segments through regular forward and backward propagations, and the shared server model segment being trained incrementally across all clients. The knowledge to be unlearned, reflected in $W_{k}^{(t)}$, would have also influenced $W_{k^{\prime}}^{(t^{\prime})}$ for all $k^{\prime}\neq k$ and $t^{\prime}\geq t$, necessitating the recalculation of corresponding changes in subsequent updates, thereby making it impossible to apply SISA unlearning [19, 12].
@@ 2.2 Machine Unlearning
Machine unlearning involves removing the influence of specific data subsets from a trained model while preserving its performance on the remaining data [10, 9]. The goal is to eliminate the impact of certain labels on the model’s predictions without compromising accuracy on the retained data. Let $D=\{\left(X_{i},Y_{q}\right)\}_{i=1,q=1}$ represent the dataset, where $X_{i}$ is the $i$-th training sample and $Y_{q}\in\{1,\cdots,Q\}$ is its label. Unlearning removes the influence of subset $D_{u}\subset D$ with labels $Y_{u}$ while preserving performance on the remaining data $D_{r}=D\backslash D_{u}$ with labels $Y_{o}$, ensuring $Y_{u}$ is forgotten and predictions on $Y_{o}$ remain accurate.
@@
Why SISA-based SL. Given the unique challenges of multi-client SL, the ability to perform machine unlearning with absent clients due to network issues is essential. While retraining and other alternatives may focus on either effectiveness or efficiency, SISA enhances these methods by enabling independent unlearning for requesting clients without affecting others. In SISA, any unlearning method can be applied within each division, converting non-SISA structures into SISA-based ones. This capability is crucial in SL, making SISA the most viable method for addressing the limitations of current non-SISA methods in SL. For simplicity, this paper assumes retraining is applied on each client.
3 SISA-based Split Unlearning
@@ 3.1 Problem Definition
In the context of SL, the training set $D_{o}$ is distributed among $K$ clients, and a server that serves as a training service provider. In the rest of this paper, we consider that the dataset is distributed in a horizontally non-IID setting of data distribution, i.e., each client owns a subset of $D_{o}$ with different feature distributions and uneven label distributions.
@@
Multi-client settings. We focus on multi-client scenarios. For a trained model $\mathcal{F}_{o}$, we can divide it into a client-side part $\mathcal{F}_{o}^{k}$ for client-$k$, and a server-side part $\mathcal{F}_{o}^{s}$. Each client-$k$ owns its $\mathcal{F}_{o}^{k}$ in parallel and shares a single $\mathcal{F}_{o}^{s}$. We define the evaluation of model performance from the perspective of client-$k$ as $\mathcal{T}(Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s}),D_{o}^{k})$, where $\mathcal{T}$ represents the evaluation function. Each client functions as an individual user managing its local data, aiming to enhance the prediction accuracy of its pre-trained local model by leveraging the server’s computational capabilities. Consequently, the performance assessment focuses on the server’s perspective, which aggregates the independent evaluations, $\mathcal{T}(Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s}),D_{o}^{k}),\forall k$.
@@
Independent unlearning. Accordingly, any client-$k$ that raises an unlearning request should be conducted only between this particular client and the server, with no needs of other $(k^{\prime}\neq k)$-th clients being involved in the unlearning process. This means the server should unlearn client-$k$’s unlearned features from the server-side part $\mathcal{F}_{o}^{s}$ upon the intermediate outputs of the updated client-side part $\mathcal{F}_{o}^{k}$. Formally, to respond to the unlearning request for a series of data samples $D_{u}^{k}$ raised by client-$k$, the entire SL network needs to satisfy an unlearned model $\mathcal{F}_{u}$ trained on $D_{u}=D_{o}\setminus\{D_{u}^{k}\}$.
@@
Challenges of SISA-based split unlearning. The split unlearning process faces what we term the Dilemma of Independence and Effectiveness, particularly when clients operate independently with limited training resources. Non-SISA methods, involving all clients, are vulnerable to client absence during unlearning and impose significant overhead on both the initiating client ($k$) and other clients ($k^{\prime}\neq k$). This method unfairly forces disinterested clients-$k^{\prime}$ to expend resources solely to meet client-$k$’s needs.
@@
Figure 1: Architectural Design. In SplitWiper (Sec.3), each client-$k$ trains its local model $\mathcal{F}_{o}^{k}$ on its dataset $D_{o}^{k}$, treating them as SISA shards without additional partitioning. After training, clients freeze the weights and send outputs with labels to the server for caching (one-off), while the server continues training $\mathcal{F}_{o}^{s}$ using cached values without returning to the clients (one-way). SplitWiper+ (Sec.4) enhances privacy by using a label expansion strategy to convert real labels into masked ones, which clients then share with the server for label-protected training.
@@
The pursuit of SISA-based split unlearning is to devise an unlearning algorithm fulfilling the following goals:
@@
    G1: Independent unlearning. This goal emphasizes the importance of performing the unlearning operation without affecting other clients ($k^{\prime}\neq k$) that are not involved in the unlearning process. This ensures that the unlearning is confined to the requesting client and server, leaving the operations and resources of other clients untouched.
@@
    G2: Effective unlearning with retained utility. This goal focuses on ensuring that the unlearning process thoroughly removes the influence of the unlearned data from the model, while also maintaining or enhancing the predictive accuracy for the retained data.

### v4

@@ Abstract
We introduce Split Unlearning, a novel machine unlearning technology designed for Split Learning (SL), enabling the first-ever implementation of Sharded, Isolated, Sliced, and Aggregated (SISA) unlearning in SL frameworks. Particularly, the tight coupling between clients and the server in existing SL frameworks results in frequent bidirectional data flows and iterative training across all clients, violating the “Isolated” principle and making them struggle to implement SISA for independent and efficient unlearning. To address this, we propose SplitWiper with a new one-way-one-off propagation scheme, which leverages the inherently “Sharded” structure of SL and decouples neural signal propagation between clients and the server, enabling effective SISA unlearning even in scenarios with absent clients. We further design SplitWiper+ to enhance client label privacy, which integrates differential privacy and label expansion strategy to defend the privacy of client labels against the server and other potential adversaries. Experiments across diverse data distributions and tasks demonstrate that SplitWiper achieves 0% accuracy for unlearned labels, and 8% better accuracy for retained labels than non-SISA unlearning in SL. Moreover, the one-way-one-off propagation maintains constant overhead, reducing computational and communication costs by 99%. SplitWiper+ preserves 90% of label privacy when sharing masked labels with the server.
@@ 1 Introduction
Split Learning (SL) is one of the latest approaches to distributed learning, suitable for scenarios where clients with limited computational resources, such as smartphones or IoT devices, need to offload part of their training to the servers [1, 2, 3]. Existing SL frameworks include two configurations: Vanilla SL and U-shaped SL [4]. In Vanilla SL, the model is split into two segments: The clients manage the segment receiving the training data and send intermediate values to the server, while the server handles the segment producing the final outputs without accessing raw data. U-shaped SL extends this by eliminating the need for label sharing, protecting both label privacy and data privacy. The model in U-shaped SL is divided into three segments: The clients process the input data and produce output classes, while the server manages the intermediate layers.
@@
The need for the compliance with regulatory requirements for “the right to forget” (e.g., GDPR [5]), which grants individuals the right to remove their personal data or impact of the data be erased, necessitates the development of effective machine unlearning [6] techniques. However, the requirement becomes challenging within the context of SL, where the model segments across clients and servers need to be jointly updated. Current methods include retraining the model [7, 8] or using alternatives like fine-tuning, gradient ascent, and knowledge distillation [9, 10, 11]. While these alternatives are favored for their higher unlearning efficiency (i.e. lower time consumption) compared to retraining, which prioritizes effectiveness (i.e., accuracy[12], they all require all data samples during each update iteration. This becomes problematic when some data samples are absent due to contextual constraints. The Sharded, Isolated, Sliced, and Aggregated (SISA) method [6, 13, 14] addresses this by segmenting data and training multiple models in parallel, maintaining higher effectiveness and efficiency than non-SISA unlearning methods.
Figure 1: Why we need SplitWiper/SplitWiper+? The “Isolated” principle of SISA invalidates existing SLs, hindering efficient and effective split unlearning. Our solution that features a novel one-way-one-off design and DP-based label extension strategy enables SISA unlearning and fulfills the fundamental requirements of SLs.
@@
Existing SL solutions [1, 2, 3, 15] are inadequate for directly applying SISA unlearning in multi-client scenarios, where multiple clients have independent models while sharing a common model on the server. Although the clients’ training data can be considered “Sharded”, SL undermines the “Isolated” requirement of SISA because: (1) the server model is tightly coupled with the client models through frequent bidirectional propagation, and (2) the shared server model is iteratively trained across all clients. Therefore, a new framework is necessary to decouple the shared server model from the client models, enabling SISA unlearning in multi-client SL environments.
@@
Contributions. In this paper, we address these gaps in enabling SISA unlearning in multi-client SL environments where unlearning can thus be conducted smoothly even with absent clients (cf. Fig. 1).
@@
2 Related Work
Unlearning in distributed environments. Machine unlearning has not been fully explored within current distributed learning frameworks, including SL. While efficient unlearning strategies have gained attention in Federated Learning (FL), which also involves collaboration between clients and a central server. Some researchers have worked on accelerating retraining across all clients, the most robust unlearning method in FL, using optimized algorithms like distributed Newton-type updates [16], SGA-EWC [17], and SFU [18]. Solutions like FedEraser [19] reduce the impact on non-unlearning-involved clients through calibration but still require cooperation from some clients. However, in SL, where client resources are often limited, non-requesting clients may lack the incentive or ability to contribute, especially if they are uninterested in or unable to participate in the unlearning process due to network issues, such as low bandwidth, high latency, or intermittent connectivity. This presents a significant challenge in SL compared to other distributed environments like FL, where developing unlearning methods that accommodate these constraints remains an unresolved issue.
SISA guideline for SL unlearning. SISA [6] is a state-of-the-art guideline for robust and efficient machine unlearning, enabling rapid adaptation to data removal requests without requiring full model retraining. By strategically partitioning data into distinct “shards” linked to different model versions, the SISA guideline allows for the targeted unlearning of specific data shards without requiring the participation of all shards. This method has been successfully applied in various domains, including unlearning in graph neural networks [13] and large language models [20]. However, SISA unlearning faces significant challenges in multi-client SL scenarios due to the interdependence between client model segments and the shared server model segment, violating the Isolated requirement of SISA. Consequently, the direct application of conventional SISA is impractical for existing SL scenarios.
Label privacy under SL unlearning. Beyond unlearning, privacy concerns regarding intermediate outputs and labels in SL have been widely studied. U-shaped SL [4] avoids label propagation to the server by placing the output layer on the client side but violates SISA’s Isolated requirement by necessitating the return of intermediate values across all clients, disabling the independent unlearning and adding overhead to resource-limited clients. If labels are shared but privacy is required, existing research primarily addresses privacy risks by directly adding noise and obfuscating labels [21, 22] or further protecting intermediate outputs transmitted to the server. Wu et al. [23] introduced a DP mechanism to perturb these transmitted outputs, while Li et al. [24] developed the Marvell method to quantify privacy leakage and optimize noise structures. However, these methods overlook the exposure of sensitive information through label quantity and semantics. Xiao et al. [25] obfuscated both labels and outputs via multiple linear combinations to reduce association with original data but still failed to conceal label quantity, potentially exposing sensitive information. Protecting the quantity of labels is critical in label-sharing SL, as revealing the exact number of labels allows adversaries to exploit supervised learning techniques that achieve higher accuracy by relying on this information, significantly increasing the risk of inference attacks and unauthorized model reconstruction.
Ours. Our design overcomes those challenges (Table 1). SplitWiper adapts SISA for multi-client SL, starting with low-effort pre-training on the clients followed by freezing their weights. This allows for one-way-one-off propagation to the server, enabling efficient unlearning without involving non-unlearning-participating clients. Additionally, SplitWiper+ enhances label privacy by expanding, shuffling, and anonymizing real labels, with intermediate outputs expanded and protected using a DP-based mechanism.
Table 1: Comparison of existing label privacy methods. Our work provides a practical privacy-preserving solution beyond existing methods by protecting label semantics, label quantity, and intermediate values while enabling effective SISA split unlearning.
Method Protection Label Semantics Label Quantity Intermediate Values Fit for SISA Split Unlearning
RRWithPrior [21]
LPSC [22]
DP-based Defense [23]
Marvell [24]
MALM [25]
U-shaped SL [4]
Ours
@@ 3 Preliminaries
We outline the key background knowledge relevant to SL and machine unlearning. By reviewing existing techniques and identifying their limitations, we set the stage for the challenges addressed in our proposed frameworks.
@@ 3.1 Split Learning
SL is a distributed machine learning technique designed to train neural networks without sharing raw data between $K$ clients, each owning datasets $D_{o}^{1},D_{o}^{2},\ldots,D_{o}^{K}$, and the server that holds no data [4, 26, 27]. A typical SL solution involves two key factors:
@@
    Topology. A network is divided into two parts, referred to as client-side and server-side. Consider the network is a sequence of layers $\{L_{1},\ldots,L_{n}\}$. The topology step involves splitting this sequence into two subsets: The client-side layers for $K$ clients are denoted as $\{\{L_{1}^{1},\ldots,L_{m}^{1}\}$, $\ldots$, $\{L_{1}^{K},\ldots,L_{m}^{K}\}\}$ and the server-side layers $\{L_{m+1},\ldots,L_{n}\}$. The clients and server initialize their respective portions of the network randomly.
@@
    Training. Each client $k$ processes its local data through the initial layers of the model, producing an intermediate output $H_{k}^{(t)}=\mathcal{F}_{o}^{k}(X_{k};W_{k}^{(t)})$ through its layers $\{L_{1}^{k},\ldots,L_{m}^{k}\}$, where $W_{k}^{(t)}$ represents the weights of client $k$ at iteration $t$. This output is sent to the server, which continues the forward propagation through its layers $\{L_{m+1},\ldots,L_{n}\}$ and updates its own weights $W_{s}^{(t)}$ based on the gradients: $W_{s}^{(t+1)}=W_{s}^{(t)}-\eta\Phi\left(\frac{\partial\mathcal{L}_{1}}{\partial W_{s}^{(t)}},\frac{\partial\mathcal{L}_{2}}{\partial W_{s}^{(t)}},\dots,\frac{\partial\mathcal{L}_{K}}{\partial W_{s}^{(t)}}\right)$, where $\frac{\partial\mathcal{L}_{k}}{\partial W_{s}^{(t)}}$ is the gradient contribution of the loss values from client $k$ and $\Phi(\cdot)$ is a customized aggregation function. The server then sends the gradients to the clients, allowing them to update their weights: $W_{k}^{(t+1)}=W_{k}^{(t)}-\eta\frac{\partial\mathcal{L}}{\partial W_{k}^{(t)}}$.
The above processes enable training without sharing raw data or model details between clients and the server. In multi-client, single-server scenarios, two primary training methods are used: synchronizing model weights between clients after each epoch or alternating training epochs between clients and the server without synchronization [4].
Incremental training process. This process has been utilized by many learning frameworks, including FL [28, 29] and SL, where a model is continuously updated by gradually incorporating new data over time. Each iteration of back-propagation during model training needs to be sent back to every client involved in the training process and builds incrementally on the previous ones, embodying the incremental nature of SL [30]. This process enables the model to dynamically learn and adapt as new data becomes available.
The challenges of applying SISA in multi-client SL scenarios stem from the incremental training process inherent in these setups. This iterative process undermines the isolated requirement of SISA due to the tight coupling between the server and client model segments through regular bidirectional propagation and the shared server model being incrementally trained across all clients. The knowledge to be unlearned, reflected in $W_{k}^{(t)}$, would have also influenced $W_{k^{\prime}}^{(t^{\prime})}$ for all $k^{\prime}\neq k$ and $t^{\prime}\geq t$, necessitating the recalculation of corresponding changes in subsequent updates, thereby making it impossible to apply SISA unlearning [31, 12].
@@ 3.2 Machine Unlearning
Machine unlearning [32, 33, 10, 9] involves removing the influence of specific data subsets from a trained model while preserving its performance on the remaining data. The goal is to eliminate the impact of certain labels on the model’s predictions without compromising accuracy on the retained data. Let $D=\{\left(X_{i},Y_{q}\right)\}_{i=1,q=1}$ represent the dataset, where $X_{i}$ is the $i$-th training sample and $Y_{q}\in\{1,\cdots,Q\}$ is its label. Unlearning removes the influence of subset $D_{u}\subset D$ with labels $Y_{u}$ while preserving performance on the remaining data $D_{r}=D\backslash D_{u}$ with labels $Y_{o}$, ensuring $Y_{u}$ is forgotten and predictions $Y_{o}$ remain accurate.
@@
Why SISA split unlearning. In multi-client SL, the ability to unlearn data despite client absence due to network issues is crucial. Existing frameworks rely on non-SISA unlearning methods, which require retraining or similar processes from the unlearning request point. While these methods prioritize either effectiveness or efficiency, SISA enables independent unlearning for requesting clients without burdening resource-limited devices in the SL network. SISA is seamlessly compatible, allowing prevalent unlearning methods to be applied individually on each client [^1] For simplicity, this paper considers retraining is used on each client (i.e., each shard)., making it the most practical solution to address limitations in current non-SISA methods.
4 SISA Split Unlearning
In this section, we define the problem and establish clear objectives for split unlearning. We present SplitWiper, a SISA-based framework designed to address existing limitations, offering a feasible and scalable solution for split unlearning.
@@ 4.1 Problem Definition
In the context of SL, the training dataset $D_{o}$ is distributed among $K$ clients and a server, functioning as a training service provider. Throughout this paper, we assume a horizontally non-IID data distribution, where each client possesses a subset of $D_{o}$ characterized by distinct feature distributions and imbalanced label distributions.
@@
Multi-client settings. We focus on multi-client scenarios. For a trained model $\mathcal{F}_{o}$, we can divide it into a client-side part $\mathcal{F}_{o}^{k}$ for client $k$, and a server-side part $\mathcal{F}_{o}^{s}$. Each client $k$ owns its $\mathcal{F}_{o}^{k}$ in parallel and shares a single $\mathcal{F}_{o}^{s}$. We define the evaluation of model performance from the perspective of client $k$ as $\mathcal{T}(Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s}),D_{o}^{k})$, where $\mathcal{T}$ represents the evaluation function. Each client functions as an individual user managing its local data, aiming to enhance the prediction accuracy of its pre-trained local model by leveraging the server’s computational capabilities with the shared server-side model segment. Consequently, the performance assessment focuses on the server’s perspective, which aggregates the independent evaluations, $\mathcal{T}(Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s}),D_{o}^{k}),\forall k$.
@@
Independent unlearning. Accordingly, any client $k$ that raises an unlearning request should be conducted only between this particular client and the server, with no needs of other clients $k^{\prime}$ $(k^{\prime}\neq k)$ being involved in the unlearning process. This means the server should unlearn the unlearned features of client $k$ from the server-side part $\mathcal{F}_{o}^{s}$ upon the intermediate outputs of the updated client-side part $\mathcal{F}_{o}^{k}$. Formally, to respond to the unlearning request for a series of data samples $D_{u}^{k}$ raised by client $k$, the entire SL network needs to satisfy an unlearned model $\mathcal{F}_{u}$ trained on $D_{u}=D_{o}\setminus\{D_{u}^{k}\}$.
@@
Challenges of SISA split unlearning. The split unlearning process faces a conflict between achieving independent unlearning for each client and maintaining overall effectiveness, which we term the Dilemma of Independence and Effectiveness, particularly when clients operate independently with limited training resources. Non-SISA methods, depending on full client participation to update shared models, are vulnerable to client absence during unlearning and impose significant overhead on both the initiating client $k$ and other clients $k^{\prime}$ ($k^{\prime}\neq k$). This method forces disinterested clients $k^{\prime}$ to expend resources solely to meet the needs of client $k$.
@@
Figure 2: Architectural Design: In SplitWiper (Sec.4), each client $k$ trains its local model $\mathcal{F}_{o}^{k}$ on its dataset $D_{o}^{k}$, treating them as SISA shards without additional partitioning. After training, clients freeze the weights and send outputs with labels to the server for caching (one-off), while the server continues training $\mathcal{F}_{o}^{s}$ using stored values without returning to the clients (one-way). SplitWiper+ (Sec.5) enhances privacy by using a label expansion strategy to convert real labels into masked ones, which clients then share with the server for label-protected training.
@@
The pursuit of SISA split unlearning is to devise an unlearning algorithm fulfilling the following goals:
@@
    G1: Independent unlearning. This goal ensures that unlearning affects only the requesting client and server, leaving other clients ($k\textquoteright\neq k$) and their resources unaffected.
@@
    G2: Effective unlearning with retained utility. This is to ensure that an unlearning process thoroughly removes the influence of the unlearned data from the model, while also maintaining the predictive accuracy for the retained data.